OpenStudy (anonymous):

Integral calculus help please. How do you use the FTC (part 1) when both a and b have x?

OpenStudy (anonymous):

you would just use it as you normally would

OpenStudy (anonymous):

treat those values as a and b

OpenStudy (anonymous):

The first fundamental theorem of calculus?

hartnn (hartnn):

you need to find the derivative of that integral ?

OpenStudy (anonymous):

That was my next question

hartnn (hartnn):

\(\Large \dfrac{d}{dx}\int \limits_{f(x)}^{g(x)}h(x)dx = H(g(x))g'(x)-H(f(x))f'(x)\)

hartnn (hartnn):

basically you will need to use the chain rule

hartnn (hartnn):

since here g(x) = x g'(x) = 1 only

hartnn (hartnn):

but f(x) is a function of x, so you need to find f'(x) f(x) = sqrt x

hartnn (hartnn):

did u get what to do ?

hartnn (hartnn):

good, if you want me to verify your final answer, tell us what u got :)

OpenStudy (anonymous):

would it not be \(h(g(x))g'(x)-h(f(x))f'(x)\)? with capital H it just makes me think it's an anti-derivative

hartnn (hartnn):

^ thats correct sorry for the confusion